OpenStudy (anonymous):

i need help Factor completely: 9x2 + 42x + 49

OpenStudy (anonymous):

a^2+2ab+b^2=(a+b)^2 (3x)^2+2*(3x*7)+(7)^2 = (3x+7)^2

OpenStudy (anonymous):

so, a^2+2ab+b^2=(a+b)^2 where: a=3x b=7